OpenStudy (anonymous):

simplify 1/36-1/x^2/1/6+1/x IMPOSSIBLE

OpenStudy (anonymous):

(x^(2)+36x-6)/(36x^(2))

OpenStudy (anonymous):

the answers are 6-x/6x .... 6x/x-6 .....x-6/6x......6x/6-x

OpenStudy (anonymous):

really? are you sure?

OpenStudy (kirbykirby):

\[\frac{1}{36}-\frac{\frac{1}{x^2}}{\frac{1}{6}}+\frac{1}{x}\]??

OpenStudy (kirbykirby):

It's really not clear how you wrote it. Use parentheses please, or LaTeX

OpenStudy (anonymous):

(1)/(36)-(1)/(x^(2))/(1)/(6)+(1)/(x) To divide by 1, multiply by the reciprocal. (1)/(36(-(1)/(x^(2))*(1)/(1)))/(6)+(1)/(x) Any expression divided by 1 is the expression. (1)/(36(-(1)/(x^(2))*1))/(6)+(1)/(x) Multiply -1 by 1 to get -1. (1)/(36(-(1)/(x^(2))))/(6)+(1)/(x) Multiply the factor by the rest of the expression to remove the fraction from the denominator. To multiply by a factor in the denominator, multiply by 1 over the factor. (1)/(36)+(1)/(6)*-(1)/(x^(2))+(1)/(x) Multiply 1 by -1 to get -1. (1)/(36)-(1)/(x^(2)*6)+(1)/(x) Multiply 6 by x^(2) to get 6x^(2). (1)/(36)-(1)/(6x^(2))+(1)/(x) To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 36x^(2). Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (1)/(36)*(x^(2))/(x^(2))-(1)/(6x^(2))*(6)/(6)+(1)/(x)*(36x)/(36x) Multiply 1 by x^(2) to get x^(2). (x^(2))/(x^(2)*36)-(1)/(6x^(2))*(6)/(6)+(1)/(x)*(36x)/(36x) Multiply 36 by x^(2) to get 36x^(2). (x^(2))/(36x^(2))-(1)/(6x^(2))*(6)/(6)+(1)/(x)*(36x)/(36x) Multiply -1 by 6 to get -6. (x^(2))/(36x^(2))-(6)/(6*6x^(2))+(1)/(x)*(36x)/(36x) Multiply 6x^(2) by 6 to get 36x^(2). (x^(2))/(36x^(2))-(6)/(36x^(2))+(1)/(x)*(36x)/(36x) Multiply 1 by 36x to get 36x. (x^(2))/(36x^(2))-(6)/(36x^(2))+(36x)/(36x*x) Multiply x by 36x to get 36x^(2). (x^(2))/(36x^(2))-(6)/(36x^(2))+(36x)/(36x^(2)) Combine the numerators of all expressions that have common denominators. (x^(2)-6+36x)/(36x^(2)) Reorder the polynomial x^(2)-6+36x alphabetically from left to right, starting with the highest order term. (x^(2)+36x-6)/(36x^(2))